if 2 is added to five times a number, the result is equal to 5 more than four times the number. Find the number.
Answer by ikleyn(52933)

.

2 + 5x = 5 + 4x


5x - 4x = 5 - 2


   x    = 3.    ANSWER

Solved.

------------

The complexity level is   1 + 2 = 3.